HanAkoSolo (jamierox4ev3r):

Even more math review! 8b. Solve the equation. (Find only the real solutions)

HanAkoSolo (jamierox4ev3r):

Posting equation in a few seconds...

HanAkoSolo (jamierox4ev3r):

\(\huge\frac{2x}{x+1} = \huge\frac{2x-1}{x}\)

Nnesha (nnesha):

any idea how to start ?

HanAkoSolo (jamierox4ev3r):

hmm... let's see. I assume you would multiply both sides by the conjugate?

OpenStudy (dinamix):

x=1

HanAkoSolo (jamierox4ev3r):

@dinamix no direct answers please. Thank you

OpenStudy (dinamix):

ok, sorry

Nnesha (nnesha):

hmm `cross multiply` \[\huge\rm \frac{ a }{ b}=\frac{ c }{ d } \]\[\huge\rm ad =bc\]

OpenStudy (nincompoop):

what happens when you do that, Jamie? There are maybe different way to approach, but let us see what will happen

OpenStudy (nincompoop):

but if we made it our goal so that we do not have to deal with denominators since they do not look the same, then cross multiplication would be an easier route

HanAkoSolo (jamierox4ev3r):

oh right. Cross multiplying is a thing. And I tried multiplying by a conjugate, but that actually seems to get kind of messy.

HanAkoSolo (jamierox4ev3r):

yep the denominators are definitely not the same

Nnesha (nnesha):

when you see `equal` sign between two fractions then u should `cross multiply`

HanAkoSolo (jamierox4ev3r):

If you cross multiply, it seems as if you get \(2x^{2} = 2x^{2}+x-1\) And thank you @Nnesha , seems like a good rule of thumb

OpenStudy (nincompoop):

now you can equate to zero and solve for x

OpenStudy (nincompoop):

I do not understand cross-multiplication. It is not a true mathematical procedure.

HanAkoSolo (jamierox4ev3r):

o-o in that case, wouldn't x be 1.

HanAkoSolo (jamierox4ev3r):

*?

Nnesha (nnesha):

yes

Nnesha (nnesha):

you can substitute x for 1 to check ur answer

HanAkoSolo (jamierox4ev3r):

a'ight then. When substituted, things look kosher (equivalent) so looks like I'm fine. But question. @nincompoop how do you know that you can squat \(2x^{2}\) to zero?

OpenStudy (nincompoop):

the equation will be set to zero by moving all of your expressions on one side and zero on the other side

OpenStudy (dinamix):

x= 1 its easy why thiss???

HanAkoSolo (jamierox4ev3r):

@dinamix I know it seems simple to just give an answer. Sadly, this does not help the user learn. So I prefer a process to be explained so I will have a strong understanding of how to do problems that are similar in the future and @nincompoop thank you ;-;

OpenStudy (nincompoop):

cross multiplication is not a true mathematical procedure tho, but it helps

HanAkoSolo (jamierox4ev3r):

it sure does. thanks, both of you

OpenStudy (nincompoop):

in cross-multiplication, we skip fundamental algebraic steps because they are supposedly understood for someone who is been doing it over and over. I can show you ALL the steps which lead to the idea of "cross-multiplication"

HanAkoSolo (jamierox4ev3r):

If you have the time to explain, feel free. I never did understand the true mathematical processes behind cross multiplication. In my freshman algebra days, we were simply told to do it because it would give us right answers.

OpenStudy (nincompoop):

it is a good way to do the review since you are about to do calculus soon, which require rigorous algebraic acrobat knowledge.

HanAkoSolo (jamierox4ev3r):

indeed. and I am quite aware. I've decided to focus on algebra the most for my review, since trig, conics, and other pre-calculus things are still fresh in my mind.

OpenStudy (dinamix):

@Jamierox4ev3rok , @nincompoop forgive me i f i was rude

OpenStudy (dinamix):

sorry anway

OpenStudy (nincompoop):